OpenStudy (anonymous):

If a rectangular prism block of wood has dimension 7 cm * 5 cm * 4 cm and cost $40 , what is the fair price in dollars of a 14 cm * 20 cm * 24 cm block of the same type if price is determined solely by volume?

OpenStudy (anonymous):

What have you attempted so far?

OpenStudy (anonymous):

i know i need to set up a ratio but I do not know how.

OpenStudy (anonymous):

Well do you know how to find the ratio of the volume of the bigger block to the smaller block?

OpenStudy (anonymous):

No :(

OpenStudy (anonymous):

For the big block you have: $$ 14 cm,~20 cm,~24 cm $$ For the small block you have: $$ 7 cm,~5 cm,~4 cm $$ Can you construct a ratio from this?

OpenStudy (anonymous):

40: 7*5*4 40:140 4:14 0.286 x:14*20*24 x:6720 0.286 = x/6720 0.286(6720) = x x = 1921

OpenStudy (anonymous):

@kevin4762 what?

OpenStudy (anonymous):

Hmm, I took a different approach to this. The ratio of the volume of the big block to the smaller block is: $$u1=2 \times 4 \times 6=48$$ Then the fair price is: $$40 \times 48 = 1920$$

OpenStudy (anonymous):

okay the ratio of price to wood is $40 for a rectangular prism of wood which has a volume of 140 units cubed. So we need to determine a fair price for a rectangular prism of wood which has a volume of 6720 units cubed. The ratio of dollars to volume is 40:140 which is .286. If we solve for x to see what price will give us a ratio of .286 when the volume is 6720, then we get 1921.

OpenStudy (anonymous):

Algorithmic's method is fine as well. There's always more than one way of solving a math problem. We both got approximately the same answer.

OpenStudy (anonymous):

I like Kevin's method. More accurate. +1

OpenStudy (anonymous):

oh i understand. thanks guys